OpenStudy (anonymous):

Formula to find a point (0,y) that is equidistant from (6,1) and (-1, -2)

OpenStudy (alfie):

Do the distance of the first point with the second one. Then do the first point of the first point with the third one. You'll get two expressions. First expression = second expression. Solve for y, and you'll get it.

OpenStudy (anonymous):

using protractor is easy

OpenStudy (anonymous):

\[d=\sqrt{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}}\] since they are equidistant \[\sqrt{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}}=\sqrt{(x_{2}-x_{3})^{2}+(y_{2}-y_{3})^{2}}\] \[\sqrt{(0-6)^{2}+(y-1)^{2}}=\sqrt{(0-(-1))^{2}+(y-(-2))^{2}}\]

OpenStudy (anonymous):

I dont understand any answer, Sorry. Im really bad at math anyway to dumb it down